# Acceleration of multi-component multiple-precision arithmetic with branch-free algorithms and SIMD vectorization

### 16 Mar 2026 v1

@@ Abstract
Multiple-precision floating-point branch-free algorithms can significantly accelerate multi-component arithmetic implemented by combining hardware-based binary64 and binary32, particularly for triple- and quadruple-precision computations. In this study, we achieved benchmark results on x86 and ARM CPU platforms to quantify the accelerations achieved in linear computations and polynomial evaluation by integrating these algorithms.
@@ 1 Introduction
Recent advances in artificial intelligence (AI), underpinned by deep learning technologies, have led to the widespread adoption of large language models and generative AI systems, permeating diverse industrial sectors and everyday human activities, including software development and coding. This trend has accelerated the shortening of mantissa length of floating-point formats employed in AI computation in pursuit of higher performance. However, in scientific and engineering computing, double-precision arithmetic represents the minimum standard, and ill-conditioned problems frequently demand even higher precision. In such cases, the effective exploitation of single instruction multiple data (SIMD) instructions and OpenMP-based parallelization techniques is crucial for harnessing the full capabilities of modern hardware. Among existing multiple-precision arithmetic frameworks, arbitrary-precision arithmetic is dominated by MPFR [15], which relies on the multiple-precision natural number arithmetic (MPN) kernel of GMP [5], and MPC [3]—an arbitrary-precision complex arithmetic library built on MPFR that achieves acceleration through the Karatsuba 3M multiplication algorithm—is widely used.
@@
By contrast, fixed-precision multi-component arithmetic based on error-free transformation (EFT) techniques, which employs combinations of multiple IEEE binary32 (24-bit mantissa and 32-bit total length) and binary64 values supported by existing hardware, is represented by quadruple-double (QD) [1] for CPUs and GQD [12]—a CUDA port of QD, though dated—for GPUs. When a relatively short mantissa length suffices, such as in double-double (DD, 106-bit), triple-double (TD, 159-bit), and quadruple-double (QD, 212-bit) arithmetic implemented by combining two to four binary64 values, computations can be executed faster than with MPFR. It has further been demonstrated by Kotakemori et al.’s Lis [7] and Hishinuma et al. [6] that basic linear computations can be accelerated by leveraging AVX, AVX2, and AVX-512, which are x86-based SIMD instruction sets. Building on an existing AVX2 implementation, we have demonstrated that basic linear computations in DD, TD, and QD precision can be accelerated on Raspberry Pi and Snapdragon platforms by exploiting Neon instructions [18].
@@
Nevertheless, DD, TD, and QD arithmetic invariably require a renormalization process at the conclusion of each operation. Particularly, for TD and QD arithmetic, this step has been identified as a bottleneck impeding performance gains through SIMD vectorization. Double-precision pair arithmetic that bypasses renormalization has been proposed as a faithfully rounded pair arithmetic scheme [11], and practical sparse matrix solver implementations that exploit its efficiency have been reported [14]. The branch-free algorithms for multiple-precision floating-point arithmetic addressed in this work [17] eliminate all conditional branches from every algorithmic step, including renormalization, and additionally reduce the operation count for TD and QD arithmetic relative to conventional algorithms. Consequently, this methodology can substantially accelerate triple- and quadruple-precision multi-component arithmetic implemented by combining hardware-based binary64 and binary32 values, and is immediately applicable to existing fixed-precision multiple-precision arithmetic implementations. As the benchmarks presented in the original paper provided insufficient comparison against existing SIMD implementations, we first quantified the speedups achievable in basic linear computations by incorporating these algorithms on x86 and ARM CPU platforms. Furthermore, as a practical application, we analyzed the effectiveness of the proposed approach when applied to a multiple-precision algebraic equation solver.
@@ 2 Branch-Free Algorithms for DD, TD, and QD Addition and Multiplication
The extension of fixed-precision floating-point numbers via EFT originated with Dekker [2]. Multiple-precision arithmetic with extended mantissa length is implemented by employing EFT primitives, represented by QuickTwoSum (Algorithms 1, 2, 3), which return both the rounded result and its rounding error using only existing floating-point operations (addition $\oplus$, subtraction $\ominus$, and multiplication $\otimes$). In the following, all variables are assumed to be representable exactly as normalized floating-point numbers.
QuickTwoSum (Algorithm 1) returns the floating-point addition result $s=a\oplus b$ and its rounding error $e$, but only under the precondition $|a|\geq|b|$; provided no overflow occurs, it maintains the relation $a+b=s+e$.
@@
TwoSum (Algorithm 2) returns the floating-point addition result $s=a\oplus b$ and its rounding error $e$ unconditionally, regardless of the relative magnitudes of $a$ and $b$; it likewise maintains the relation $a+b=s+e$.
@@
TwoProd (Algorithm 3) returns the floating-point multiplication result $p=a\otimes b$ and its rounding error $e$, maintaining the relation $ab=p+e$. As virtually all floating-point hardware currently in operation supports fused multiply-add (FMA) instructions, the FMA-based variant of TwoProd was employed throughout this work.
@@
As these EFT operations are constructed entirely from existing floating-point operations and contain no conditional branches, they can be directly implemented using SIMD instructions. With AVX2, which employs 256-bit registers on x86 architectures, four binary64 or eight binary32 values can be processed simultaneously by EFT operations; with Neon on AArch64, which employs 128-bit registers, two binary64 or four binary32 values can be processed simultaneously. Computations that contain no conditional branches and can be executed without overflowing SIMD registers in this manner are termed branch-free (BF).
By combining EFT operations, multiple-precision arithmetic with an extended mantissa can be implemented by composing multiple floating-point numbers; this approach is referred to as the multi-component (or multiterm) method. Because subtraction reduces to sign negation and division can be implemented via Newton’s method combining addition and multiplication, these two operations are crucial from a performance perspective. The correct handling of rounding errors is critical in this implementation and requires a carefully derived, rigorous error analysis. As the mantissa length increases, the operation count grows, and error management becomes increasingly complex. Zhang & Aiken proposed an abstract model for automated reasoning that simplifies floating-point arithmetic [16] and introduced branch-free multi-component multiple-precision addition and multiplication algorithms. In the following, we present and discuss these BF algorithms.
@@ 2.1 Remarks on double-word (DW) Arithmetic
The simplest form of multi-component multiple-precision floating-point arithmetic is DW arithmetic. When binary64 (double-precision) is used as the base format, it is abbreviated as DD; when binary32 (single-precision) is used, it is abbreviated as DS. Algorithm 4 presents the default DW addition currently defined in QD [1], which is the sloppy or fast variant; its final rounding error is known to be larger than that of the accurate version (Algorithm 5).
@@
$(s,e):=\mbox{TwoSum}(a[0],b[0])$;
$e:=e\oplus(a[1]\oplus b[1])$;
$(c[0],c[1]):=\mbox{QuickTwoSum}(s,e)$;
@@
$(s_{1},s_{2}):=\mbox{TwoSum}(a[0],b[0])$;
$(t_{1},t_{2}):=\mbox{TwoSum}(a[1],b[1])$;
$s_{2}:=s_{2}\oplus t_{1}$;
$(s_{1},s_{2}):=\mbox{QuickTwoSum}(s1,s2)$;
$s_{2}:=s_{2}\oplus t_{2}$;
$(c[0],c[1]):=\mbox{QuickTwoSum}(s1,s2)$;
@@
Zhang & Aiken proposed Algorithm 6 as a branch-free reformulation of the accurate DW addition.
@@
$(p_{1},p_{2}):=\mbox{TwoProd}(a[0],b[0])$;
$p_{2}:=p_{2}\oplus(a[0]\otimes b[1]\oplus a[1]\otimes b[0])$;
@@
The corresponding BF multiplication is proposed as presented in Algorithm 8. This algorithm can be interpreted as performing all necessary multiplications, followed by arranging error-handling steps in a branch-free manner.
@@
Consequently, the BF variants of DW addition and multiplication involve no fewer floating-point operations than their conventional counterparts, and—as will be discussed subsequently—no measurable performance benefit was observed under the current default DW arithmetic. Therefore, in the benchmark tests reported as follows, DD arithmetic employs the conventional (nonBF) implementations unless otherwise stated; only TD and QD arithmetic utilize the BF algorithms presented hereafter.
@@ 2.2 Remarks on TW Arithmetic
For triple-precision arithmetic (Triple-Word, TW), one approach is to consider the fourth component of QD arithmetic as zero; alternatively, Fabiano et al. [4] proposed a formulation with a reduced operation count. Through SIMD implementation, we validated that the former approach outperforms the latter [9]; therefore, Algorithm 9 is used for TW addition and Algorithm 10 for TW multiplication.
@@
Both algorithms conclude with a call to the TW renormalization function TWRenormalize, which requires up to five QuickTwoSum operations and necessitates conditional branching via if-statements. This function has not yet been SIMD-vectorized and is considered a primary source of performance degradation.
In contrast, the proposed BF algorithms, presented by Algorithms 11 and 12, introduce no conditional branches in the renormalization-equivalent steps and require fewer operations than their conventional counterparts. This renders them highly amenable to performance improvement through SIMD vectorization.
@@ 2.3 Remarks on QW Arithmetic
Quadruple-precision arithmetic (Quadruple-Word, QW) is based on the implementation provided in QD; as noted previously, TD arithmetic is an abbreviated form of QD arithmetic with the fourth component zeroed. Both QD addition and QD multiplication conclude with a call to the QW renormalization function (QWRenormalize), which contains conditional branches and requires up to seven QuickTwoSum function calls.
In contrast, the BF algorithms are presented by Algorithms 13 and 14. Although the time required to discover these BF algorithms is unknown, the resulting algorithms are remarkably concise and involve a substantially reduced operation count. Consequently, the benefit of SIMD vectorization is expected to be even more pronounced for QW arithmetic than for TW arithmetic.
@@ 3 Benchmark Tests for Real and Complex Square Matrix Multiplication
Lenovo IdeaPad Slim 5(Snapdragon X Plus X1P-42-100, 3.4 GHz, 34 MB L3 cache, 8 cores)Windows 11, WSL2 + Ubuntu 24.04.2 LTS, GCC 13.3.0
Standard intrinsic functions provided as C functions were used for SIMD implementation in both environments. The Arm Neon vectorization and OpenMP parallelization were implemented primarily with the assistance of Anthropic Claude, by converting the existing AVX2 code [18].
@@ 3.1 Real Matrix Multiplication
The square matrices employed in the tests are defined as
@@
Our library supports four matrix multiplication schemes: the naive triple-loop method, a block-based algorithm, Strassen’s algorithm, and the Ozaki scheme. In the present evaluation, Strassen’s algorithm was employed to isolate the performance contribution of BF arithmetic. The recursion in Strassen’s algorithm terminates when the submatrix size falls below $32\times 32$. The performance benefit of BF arithmetic was measured for both nonvectorized and SIMD-vectorized (AVX2 and Neon) implementations. Execution times with OpenMP parallelization using 8 and 32 threads are also reported. Because our Strassen routine does not achieve further speedup beyond 8 threads, results for the block-based algorithm (B32 Threads) are additionally included for the 32-thread case only. The numerical accuracy of all computed results falls within the ranges of 29.0–30.4 significant decimal digits for DD, 45.5–47.5 for TD, and 61.7–63.5 for QD precision.
@@
Figure 1: Computation time (s) of Strassen matrix multiplication: Snapdragon (left), EPYC (right)
Notably, BF arithmetic provides no benefit—rather degrades performance—for DD precision, whereas it yields clear speedups for TD and QD precisions. For $n=2049$, the resulting speedups over the MPBLAS [13] DD- and QD-precision Rgemm are summarized in Table 1.
@@
Table 2 reports the average speedup ratios owing to BF arithmetic for DD precision, where BF arithmetic offers no benefit, comparing nonSIMD (NS) and SIMD-vectorized implementations.

### 07 May 2026 v2

# Acceleration of multicomponent multiple-precision arithmetic with branch-free algorithms and SIMD vectorization

@@ Abstract
Multiple-precision floating-point branch-free algorithms can significantly accelerate multicomponent arithmetic implemented by combining hardware-based binary64 and binary32, particularly for triple- and quadruple-precision computations. In this study, we achieved benchmark results on x86 and ARM CPU platforms to quantify the accelerations achieved in linear computations and polynomial evaluations by integrating these algorithms.
@@ 1 Introduction
Recent advances in artificial intelligence (AI), underpinned by deep-learning technologies, have led to the widespread adoption of large language models and generative AI systems, permeating diverse industrial sectors and everyday human activities, including software development and coding. This trend has accelerated the shortening of the mantissa lengths in floating-point formats employed in AI computations to achieve higher performance. However, in scientific and engineering computing, double-precision arithmetic is the minimum standard, and ill-conditioned problems frequently require higher precision. In such cases, the effective use of single-instruction multiple-data (SIMD) instructions and OpenMP-based parallelization techniques is crucial for harnessing the full capabilities of modern hardware. Among the existing multiple-precision arithmetic frameworks, arbitrary-precision arithmetic is dominated by MPFR [15], which relies on the multiple-precision natural-number arithmetic (MPN) kernel of GMP [5]. MPC [3], an arbitrary precision complex arithmetic library built on MPFR that achieves acceleration through the Karatsuba 3M multiplication algorithm, is widely used.
@@
In contrast, fixed-precision multicomponent arithmetic based on error-free transformation (EFT) techniques, which employs combinations of multiple IEEE binary32 (24-bit mantissa and 32-bit total length) and binary64 values supported by existing hardware, is represented by quadruple-double (QD) [1] for CPUs and GQD [12]—a CUDA port of QD, though dated—for GPUs. When a relatively short mantissa length suffices, such as in double-double (DD, 106-bit), triple-double (TD, 159-bit), and quadruple-double (QD, 212-bit) arithmetic implemented by combining two to four binary64 values, computations can be executed faster than with MPFR. Kotakemori et al.’s Lis [7] and Hishinuma et al. [6] further demonstrated that basic linear computations can be accelerated by leveraging AVX, AVX2, and AVX-512, which are x86-based SIMD instruction sets. Building on an existing AVX2 implementation, we demonstrated that basic linear computations in DD, TD, and QD precision can be accelerated on the Raspberry Pi and Snapdragon platforms by exploiting the Neon instructions [18].
@@
Nevertheless, DD, TD, and QD arithmetic require renormalization at the conclusion of each operation. In particular, for TD and QD arithmetic, this step has been identified as a bottleneck that impedes performance gains through SIMD vectorization. Double-precision pair arithmetic, which bypasses renormalization, has been proposed as a faithfully rounded pair-arithmetic scheme [11], and practical sparse matrix solver implementations that exploit this efficiency have been reported [14]. The branch-free algorithms for multiple-precision floating-point arithmetic addressed in this work [17] eliminate all conditional branches from every algorithmic step, including renormalization, and reduce the operation count for TD and QD arithmetic relative to conventional algorithms. Consequently, this methodology can substantially accelerate triple- and quadruple-precision multicomponent arithmetic implemented by combining hardware-based binary64 and binary32 values and is immediately applicable to existing fixed-precision multiple-precision arithmetic implementations. Because the benchmarks presented in the original paper provided an insufficient comparison against existing SIMD implementations, we first quantified the speedups achievable in basic linear computations by incorporating these algorithms on the x86 and ARM CPU platforms. Furthermore, as a practical application, we analyzed the effectiveness of the proposed approach by applying it to a multiple-precision algebraic equation solver.
@@ 2 Branch-Free Algorithms for DD, TD, and QD Addition and Multiplication
The extension of fixed-precision floating-point numbers via EFT was pioneered by Dekker [2]. Multiple-precision arithmetic with an extended mantissa length was implemented by employing EFT primitives, represented by QuickTwoSum (Algorithms 1, 2, 3), which return both a rounded result and its rounding error using only existing floating-point operations (addition $\oplus$, subtraction $\ominus$, and multiplication $\otimes$). In the following, all variables are assumed to be exactly representative as normalized floating-point numbers.
QuickTwoSum (Algorithm 1) returns the floating-point addition result $s=a\oplus b$ and its rounding error $e$, but only under the precondition $|a|\geq|b|$; if no overflow occurs, it maintains the relation $a+b=s+e$.
@@
TwoSum (Algorithm 2) returns the floating-point addition result $s=a\oplus b$ and its rounding error $e$ unconditionally, regardless of the relative magnitudes of $a$ and $b$, while maintaining the relationship $a+b=s+e$.
@@
TwoProd (Algorithm 3) returns the floating-point multiplication result $p=a\otimes b$ and its rounding error $e$ while maintaining the relationship $ab=p+e$. Because virtually all floating-point hardware currently in operation supports fused multiply–add (FMA) instructions, the FMA-based variant of TwoProd was employed throughout this work.
@@
As these EFT operations are constructed entirely from existing floating-point operations and contain no conditional branches, they can be implemented directly using SIMD instructions. With AVX2, which employs 256-bit registers on x86 architectures, four binary64 or eight binary32 values can be processed simultaneously by EFT operations. In contrast, with Neon on AArch64, which employs 128-bit registers, two binary64 or four binary32 values can be processed simultaneously. Computations that contain no conditional branches and can be executed without overflowing SIMD registers in this manner are termed branch-free (BF).
By combining EFT operations, multiple-precision arithmetic with extended mantissa can be implemented by composing multiple floating-point numbers; this is referred to as the multicomponent (or multiterm) method. Because subtraction reduces to sign negation and division can be implemented via Newton’s method that combines addition and multiplication, these two operations are crucial from a performance perspective. Correct handling of the rounding errors is critical in this implementation and requires a carefully derived rigorous error analysis. As the mantissa length increases, the operation count increases, and error management becomes increasingly complex. Zhang & Aiken proposed an abstract model for automated reasoning that simplifies floating-point arithmetic [16] and introduced branch-free multicomponent multiple-precision addition and multiplication algorithms. In the following section, we present and discuss BF algorithms.
@@ 2.1 Remarks on double-word (DW) Arithmetic
The simplest form of multicomponent multiple-precision floating-point arithmetic is the DW arithmetic. When binary64 (double-precision) is used as the base format, it is abbreviated as DD; when binary32 (single-precision) is used, it is abbreviated as DS. Algorithm 4 presents the default DW addition defined in QD [1], which is the sloppy or fast variant; its final rounding error is larger than that of the accurate version (Algorithm 5).
@@
$(s,e):=\mbox{TwoSum}(a[0],b[0])$; $e:=e\oplus(a[1]\oplus b[1])$; $(c[0],c[1]):=\mbox{QuickTwoSum}(s,e)$;
@@
$(s_{1},s_{2}):=\mbox{TwoSum}(a[0],b[0])$; $(t_{1},t_{2}):=\mbox{TwoSum}(a[1],b[1])$;
$s_{2}:=s_{2}\oplus t_{1}$; $(s_{1},s_{2}):=\mbox{QuickTwoSum}(s1,s2)$;
$s_{2}:=s_{2}\oplus t_{2}$; $(c[0],c[1]):=\mbox{QuickTwoSum}(s1,s2)$;
@@
Zhang and Aiken proposed Algorithm 6 as a branch-free reformulation for accurate DW addition.
@@
$(p_{1},p_{2}):=\mbox{TwoProd}(a[0],b[0])$; $p_{2}:=p_{2}\oplus(a[0]\otimes b[1]\oplus a[1]\otimes b[0])$;
@@
The corresponding BF multiplication is presented as Algorithm 8. This algorithm can be interpreted as performing all necessary multiplications, followed by arranging the error-handling steps in a branch-free manner.
@@
Consequently, the BF variants of DW addition and multiplication involve no fewer floating-point operations than their conventional counterparts. In addition, as discussed in the subsequent sections, they exhibit no measurable performance benefit under the current default DW arithmetic. Therefore, in the following benchmark tests, DD arithmetic employs conventional (non-BF) implementations unless otherwise stated; only TD and QD arithmetic utilize the BF algorithms presented hereafter.
@@ 2.2 Remarks on TW Arithmetic
For triple-precision arithmetic (triple-word, TW), one approach is to consider the fourth component of the QD arithmetic as zero. In contrast, Fabiano et al. [4] proposed a formulation with a reduced operation count. Through SIMD implementation, we validated that the former approach outperforms the latter [9]; therefore, Algorithm 9 is used for TW addition and Algorithm 10 for TW multiplication.
@@
Both algorithms conclude with a call to the TW renormalization function TWRenormalize[4][1], which requires up to five QuickTwoSum operations and necessitates conditional branching via if-statements. This function has not yet been SIMD-vectorized and is considered the primary source of performance degradation.
By contrast, the proposed BF algorithms presented in Algorithms 11 and 12 introduce no conditional branches in the renormalization-equivalent steps and require fewer operations than their conventional counterparts. This renders them highly amenable to performance improvements through SIMD vectorization.
@@
As listed in Table 1, the branch-free triple-word multiplication has the least total number of primary computations compared with our currently adopted and Fabiano’s and triple-word multiplications. However, we could not find a clear difference among the three types of triple-word additions.
Table 1: Operation counts for Triple-Word arithmetic
Triple-Word (TW) Add Sub Mul Div FMA
TW+TW 19 34 0 0 0
Fabiano TW+TW 20 40 0 0 0
BF TW+TW 21 36 0 0 0
[…]
@@ 2.3 Remarks on QW Arithmetic
Quadruple-precision arithmetic (quadruple-word, QW) is based on the implementation provided in QD. As noted previously, TD arithmetic is an abbreviated form of QD arithmetic with the fourth component zeroed. Both the QD addition and multiplication conclude with a call to the QW renormalization function (QWRenormalize[1]), which contains conditional branches and requires up to seven QuickTwoSum function calls.
The BF algorithms are presented in Algorithms 13 and 14. Although the time required to discover these BF algorithms is unknown, the resulting algorithms are remarkably concise and involve substantially reduced operation counts. Consequently, the benefit of SIMD vectorization is expected to be more pronounced for QW arithmetic than for TW arithmetic.
@@
We present Table 2 for comparison with our adopted and branch-free quadruple-word algorithms. Similarly, as listed in Table 1, branch-free multiplication is the most lightweight, and no significant difference exists between the two types of quadruple-word additions.
Table 2: Operation counts for quad-word arithmetic
Quad-Word (QW) Add Sub Mul Div FMA
QW+QW 26 58 0 0 0
BF QW+QW 37 66 0 0 0
[…]
@@ 3 Benchmark Tests for Real and Complex Square Matrix Multiplication
Lenovo IdeaPad Slim 5 (Snapdragon X Plus X1P-42-100, 3.4 GHz, 34 MB L3 cache, 8 cores), Windows 11, WSL2 + Ubuntu 24.04.2 LTS, GCC 13.3.0
Standard intrinsic functions, provided as C functions, were used for SIMD implementation in both environments. Arm Neon vectorization and OpenMP parallelization were implemented primarily with assistance from the Anthropic Claude by converting the existing AVX2 code [18].
@@ 3.1 Real Matrix Multiplication
The square matrices employed in the tests are defined as follows:
@@
Our library supports four matrix multiplication schemes: the naive triple-loop method, a block-based algorithm, Strassen’s algorithm, and the Ozaki scheme. In this study, Strassen’s algorithm was employed to isolate the performance contribution of the BF arithmetic. The recursion in Strassen’s algorithm terminates when the submatrix size decreases below $32\times 32$. The performance benefit of BF arithmetic was measured for both nonvectorized and SIMD-vectorized (AVX2 and Neon) implementations. Execution times with OpenMP parallelization using 8 and 32 threads were also reported. Because our Strassen routine did not achieve further speedup beyond eight threads, the results for the block-based algorithm (B32 threads) were included only for the 32-thread case. The numerical accuracy of all computed results falls within the ranges of 29.0–30.4 significant decimal digits for DD, 45.5–47.5 for TD, and 61.7–63.5 for QD precision.
@@
Figure 1: Computation time (s) of real-valued Strassen matrix multiplication: Snapdragon (left), EPYC (right)
Notably, BF arithmetic provides no benefit and degrades performance for DD precision, whereas it yields clear speedups for TD and QD precisions. The resulting speedups over the MPBLAS [13] DD- and QD-precision Rgemm for $n=2049$ are summarized in Table 3.
@@
Table 4 lists the average speedup ratios owing to BF arithmetic for DD precision, where BF arithmetic offers no benefit compared to nonSIMD (NS) and SIMD-vectorized implementations.